### v1

# Countering adversarial evasion in regression analysis

Abstract
Adversarial machine learning challenges the assumption that the underlying distribution remains consistent throughout the training and implementation of a prediction model. In particular, adversarial evasion considers scenarios where adversaries adapt their data to influence particular outcomes from established prediction models, such scenarios arise in applications such as spam email filtering, malware detection and fake-image generation, where security methods must be actively updated to keep up with the ever-improving generation of malicious data. Game theoretic models have been shown to be effective at modelling these scenarios and hence training resilient predictors against such adversaries. Recent advancements in the use of pessimistic bilevel optimsiation which remove assumptions about the convexity and uniqueness of the adversary’s optimal strategy have proved to be particularly effective at mitigating threats to classifiers due to its ability to capture the antagonistic nature of the adversary. However, this formulation has not yet been adapted to regression scenarios. This article serves to propose a pessimistic bilevel optimisation program for regression scenarios which makes no assumptions on the convexity or uniqueness of the adversary’s solutions.
1 Introduction
Adversarial machine learning considers the exploitable vulnerabilities of machine learning models and the strategies needed to counter or mitigate such threats [32]. By considering these vulnerabilities during the development stage of our machine learning models, we can work to build resilient methods [9, 11] such as protection from credit card fraud [35] or finding the optimal placement of air defence systems [20]. In particular, we consider the model’s sensitivity to changes in the distribution of the data. The way the adversary influences the distribution can fall under numerous categories, see [21] for a helpful taxonomy that categorises these attacks. We focus on the specific case of exploratory attacks, which consider the scenarios where adversaries attempt to modify their data to evade detection by a classifier. Such attacks might occur in security scenarios such as malware detection [3] and network intrusion traffic [31]. In a similar vein, and more recently, vulnerabilities in deep neural networks (DNN) are being discovered, particularly in the field of computer vision and image classification; small perturbations in the data can lead to incorrect classifications by the DNN [33, 19]. These vulnerabilities raise concerns about the robustness of the machine learning technology that is being adopted and, in some cases, in how safe relying on their predictions could be in high-risk scenarios such as autonomous driving [15] and medical diagnosis [16]. By modelling the adversary’s behaviour and anticipating these attacks, we can train classifiers that are resilient to such changes in the distribution before they occur.
Game theory provides an effective and popular technique to model adversarial scenarios. These games see one player, the learner, attempt to train a prediction model while another player, the adversary, modifies their data in an attempt to influence particular outcomes from the learner’s predictor. See, for example [10] for early work in modelling adversarial learning. The precise structure of such a model then depends on a number of factors, such as whether the attack occurs either at training time [27], or implementation time [5]. Further to this, while some games assume the players act simultaneously [4, 29], others allow for sequential play where either the adversary acts first [25, 26, 7, 24] or the learner acts first [21, 5]. We focus our attention on attacks made at implementation time where the adversary seeks to influence an already established prediction model, commonly known as evasion attacks in the literature [21]. Sequential evasion games naturally take on the form of a bilevel optimisation program. In this formulation, the learner, taking the role of the leader in the upper-level, train a prediction function while anticipating how the adversary, taking the role of the follower in the lower-level, will react. The solution to this program then results in a resilient prediction function. Pessimistic bilevel optimisation in particular has proved particularly effective. When multiple optimal solutions are available to the adversary, the pessimistic model, unlike its optimistic variant, assumes no cooperation between the adversary and the learner. This formulation better captures the antagonistic nature of adversarial evasion scenarios [2, 5].
In this article, we adapt the pessimistic bilevel program proposed by [37] to train resilient predictors in regression scenarios. We model scenarios where we might expect adversaries to modify their data in an attempt to influence a particular label from a prediction model. We use the motivating example of an adversary providing falsified information about a house in order to influence a higher valuation. We explore properties of the constrained bilevel model under new choices of objective functions, showing the possible existence of multiple optimal solutions and hence the benefit of a constrained pessimistic bilevel approach, which, to the best of our knowledge, has not yet been applied to regression analysis. Moreover, we take this opportunity to exploit the fact that our method does not require the use of feature maps in its solution method. Feature maps, such as PCA, as was used [5], are typically not bijective. Consequently, although a solution to the adversary can be found in the map-induced feature space, a corresponding solution in the original feature space cannot be identified. However, under our model, we can view the explicit values of the adversary’s data at their solution. These values can give us insight into the strategies played by the adversary and allows us to identify the extent to which each feature needs to be modified in order to adequately influence the prediction model. From this, we can identify which features are most vulnerable. For example, if a certain feature required only a minor perturbation, then we might consider this feature to be particularly vulnerable to adversarial influence. On the other hand, if a feature required large transformations to trick the learner, then we might consider this feature to be resilient to attacks.
The contributions of this article can be summarised as follows. We adapt the bilevel model proposed in [37] to accommodate regression analysis and present a novel pessimistic bilevel optimisation program to train prediction functions that are resilient to strategic changes in the data of regression scenarios. This bilevel program makes no assumptions on the convexity of the lower-level problem, or the uniqueness of its solution, which, to the best of our knowledge, has not previously been proposed for adversarial regression analysis. We design experiments to assess the performance of our model and showcase its ability to outperform existing methods. Finally, we demonstrate a key benefit of our model over existing approaches. Where existing approaches required the use of kernel transformations to ensure convexity, we are able to keep the lower-level variable (the adversary’s data) in its original feature space, allowing us to investigate the value of the data at the solution. From this we can infer which features require the greatest transformation in order to effectively influence the learner’s prediction function.
The remainder of this paper is outlined as follows; we begin in Section 2 by presenting the constrained pessimistic bilevel model to train resilient prediction functions in regression analysis before outlining the solution method. Following this, we assess the performance of the bilevel model on two datasets in Section 3 before concluding our findings in Section 4.
2 Training resilient prediction functions
In this section we present an adaptation of the constrained pessimistic bilevel program proposed in [37] to model adversarial regression. In particular, we demonstrate how the process of adversarial training can be used to train resilient predictors under the scenario of adversaries attempting to strategically modify their data in an attempt to be assigned incorrect labels from established prediction models. Such scenarios might arise, for example, in property valuation. An adversary might lie and provide falsified information in an attempt to influence a higher valuation. Through suitable choices of objective functions, we construct a constrained pessimistic bilevel program that sees a learner, in the upper-level, train a prediction function while considering how an adversary, in the lower-level, might alter their data towards a target outcome from the prediction model. The adversary is assigned some existing data as a starting point. The constraints are then used to measure the similarity between the adversary’s solution and starting point. As demonstrated in [37], these constraints ensure that the adversary’s data remain plausibly realistic and prevent instances of data losing their intended meaning. In this way, the bilevel model more realistically simulates adversarial transformations, leading to improved performance. To the best of our knowledge, a constrained pessimistic bilevel program with no assumptions on convexity or uniqueness of lower-level solution has previously been proposed for adversarial evasion scenarios in regression analysis.
Let $x\in\mathbb{R}^{q}$ be a sample (row) of data containing $p\in\mathbb{N}$ features with corresponding label $y\in\mathbb{R}$. For some weights $w\in\mathbb{R}^{q}$, we define the learner’s prediction function, $\sigma:\mathbb{R}\times\mathbb{R}\rightarrow\mathbb{R}$, as the linear weighted sum of the features,
$\sigma(w,x):=w_{0}+w^{T}x.$ (1)
This prediction function is a common choice, for example, in linear regression. We then construct a loss function, $\mathcal{L}:\mathbb{R}\times\mathbb{R}\rightarrow\mathbb{R}$, to train the prediction function and identify appropriate weights. As in linear regression, we can achieve this by minimising the squared error between the learner’s prediction of $x$ and the label, $y$,
$\mathcal{L}(\sigma(w,x),y):=(\sigma(w,x)-y)^{2}=(w_{0}+w^{T}x-y)^{2}.$
The adversary, in the lower-level, transforms some data with the aim of having the learner’s prediction function, $\sigma$ output a desired target label. For example, consider again the example of an adversary lying about features of a housing property to influence a higher valuation than deserved. If, for example, the actual valuation price is $y\in\mathbb{R}$, the adversary might set their target, $z\in\mathbb{R}$, as $z=y+\nu$, where $\nu>0$. The lower-level (adversary’s) loss function is then defined as the squared distance between the learner’s prediction and the target label,
$\ell(\sigma(w,x),z):=(\sigma(w,x)-z)^{2}=(w^{T}x-z)^{2}.$ (2)
With the loss functions established, we now work them into objective functions that accommodate a full dataset. We divide the training data into two sets, a static set as would typically be used in the training of a machine learning predictor, and then second set of data which can be manipulated by the adversary. Let $D\in\mathbb{R}^{nq},D=(D_{1},\dots,D_{n})^{T}$ be the static set of $n\in\mathbb{N}$ instances of data where each $D_{i}\in\mathbb{R}^{q},\;i=1,\dots,n$, is a row vector containing the values of $q\in\mathbb{N}$ features, and let $\gamma\in\{0,1\}^{n}$ be the corresponding corresponding collection of labels. Let $X\in\mathbb{R}^{mq}$ be the set of $m\in\mathbb{N}$ instances which can be manipulated by the adversary, defined as,
$X:=\begin{pmatrix}X_{1}^{T}\\ \vdots\\ X_{m}^{T}\end{pmatrix}$,
where each $X_{i},\;i=1,\dots,m$, is a row vector of features. Let $Y\in\{0,1\}^{m}$ be the corresponding labels of the adversary’s data. The upper level objective of the bilevel program sees the learner minimising their loss over both sets of data. Note that the learner considers the true labels of the adversary’s data, where $F:\mathbb{R}^{p}\times\mathbb{R}^{mp}\rightarrow\mathbb{R}$ is the upper-level objective,
$F(w,X):=\frac{1}{n}\sum_{i=1}^{n}\mathcal{L}(\sigma(w,D_{i}),\gamma_{i})+\frac{1}{m}\sum_{i=1}^{m}\mathcal{L}(\sigma(w,X_{i}),Y_{i}).$
Then, given a set of weights, $w$ of the learner’s prediction function, the adversary, in the lower-level seeks to modify their data such that the learner’s prediction moves towards their target labels, $Z$. Let $f:\mathbb{R}^{p}\times\mathbb{R}^{mp}\rightarrow\mathbb{R}$ be the adversary’s objective, which is defined as the sum of the adversary’s loss over their data,
$f(w,X):=\frac{1}{m}\sum_{i=1}^{m}\ell(\sigma(w,X_{i}),Z_{i}).$ (3)
Finally, we define the constraints $g:\mathbb{R}^{mq}\rightarrow\mathbb{R}^{m}$ as
$g(X):=\begin{pmatrix}g_{1}(X_{1})\\ \vdots\\ g_{m}(X_{m})\end{pmatrix}$, (4)
where each constraint function, $g_{i}:\mathbb{R}^{q}\rightarrow\mathbb{R},\;i=1,\dots,m$, is defined as
$g(X):=\delta-d(X_{i},X^{0}_{i}),$ (5)
where $d:\mathbb{R}^{q}\times\mathbb{R}^{q}\rightarrow\mathbb{R}$ is some similarity function. These constraints serve to measure and restrict the cosine similarity between the adversary’s solution and initial value to be below some threshold $\delta\in\mathbb{R}$. As outlined in [37], these constraints prevent the adversary’s data becoming nonsensical and too far removed from its intended message. This continues to be useful under the regression scenarios as we can reduce the likelihood of the adversary producing unrealistic data. For example, in reality, an adversary cannot claim that their house was built thousands of years ago when in reality it was built a mere 5 years ago. This would be immediately suspicious to anybody who checks the adversary’s data. However, smaller transformations, such as within the range of $\pm 2$ years would be far more believable. Similarly to the work in [37], we use the cosine similarity to measure the extent to which the adversary has modified their data. Let $X^{0}\in\mathbb{R}^{mq}$ be the initial value of the adversary’s data, then the constraints are then defined as follows,
$g_{i}\left(X\right)=\delta-d(X_{i},X^{0}_{i})=\delta-\frac{X_{i}\cdot X_{i}^{0}}{\|X_{i}\|\|X_{i}^{0}\|},\;i\in\{1,\dots,m\}.$ (6)
Then, setting the initial value of the adversary’s data to some existing data in the training set, we can be confident that under these constraints, the adversary produces data that are similar in nature to some real example.
We now formulate the interactions between the learner and adversary as a bilevel optimisation problem. To begin, given a set of weights, $w$, of the learner’s prediction function, let $S(w)$ be the set of optimal data for the adversary,
$S(w):=\operatorname*{argmin}_{X\in\mathbb{R}^{mq}}\left\{f(w,X)\mid g(X_{i})\leq 0\right\},$ (7)
This set contains all instances of data that will trick the learner into assigning the adversary their desired labels $Z$. For example, the falsified falsified report that gives a property a better valuation than it actually deserves. Note that the adversary’s objective function under the linear regression loss in (2) and prediction function (1) possess multiple optimal solutions. This can be directly shown using in [37, Proposition 1] since the prediction function, $\sigma$, can be expressed as a function of the linear combination $w_{0}+w^{T}x$ for some $x\in\mathbb{R}^{q}$. Simply let $\Omega(w_{0}+w^{T}x)=w_{0}+w^{T}x$ and the proof follows.
Note also that its has already demonstrated in [37, Proposition 2] that the feasible region defined by these constraints is non-convex. Therefore, in the case of multiple optimal solutions to the adversary’s problem, given by the set $S(w)$, we make the pessimistic assumption that the adversary will not act cooperatively, namely, they will choose the solution which maximises the learner’s objective, while the learner seeks to minimise it. The complete bilevel program is then given by
$\min_{w\in\mathbb{R}^{p}}\max_{X\in S(w)}F(w,X).$ (8)
The non-convexity and existence of multiple optimal solutions to the lower-level problem motives the use of the solution method proposed in [37]. To begin, we collect the upper and lower variables, $w$ and $X$, into the block variable $z$ as below
$z:=\begin{bmatrix}w\\ X\end{bmatrix}\in\mathbb{R}^{q+mq},\qquad$
and define the block variable $\zeta\in\mathbb{R}^{2b+1}$ as,
$\xi:=\begin{bmatrix}\beta\\ \hat{\beta}\\ \lambda\end{bmatrix}\in\mathbb{R}^{2b+1},$
where $\beta,\hat{\beta}\in\mathbb{R}^{b}$ and $\lambda\in\mathbb{R}$ are lagrange multiples with associated Lagrangian-type functions $L^{\textup{p}}_{w},L^{\textup{p}}_{X},\ell^{\textup{p}}\colon(\mathbb{R}^{q}\times\mathbb{R}^{mq})\times\mathbb{R}^{2b+1}\to\mathbb{R}$ as stated below:
$L^{\textup{p}}_{w}(z,\xi):=-F(w,X)+(\beta-\lambda\hat{\beta})^{\top}g(X)$
$L^{\textup{p}}_{X}(z,\xi):=-F(w,X)+\lambda f(w,X)+\beta^{\top}g(X)$
$\ell^{\textup{p}}(z,\xi):=f(w,X)+\hat{\beta}^{\top}g(X).$
Now, we can define the block functions, $\mathcal{G}$ and $\mathcal{H}$ as below,
$\mathcal{G}(z):=\begin{bmatrix}g(X)\\ g(X)\\ 0\end{bmatrix},\qquad\mathcal{H}(z,\xi;B_{i}):=\begin{bmatrix}\nabla_{w}L^{\textup{p}}_{w}(z,\xi)\\ \nabla_{X}L^{\textup{p}}_{X}(z,\xi)\\ \nabla_{X}\ell^{\textup{p}}(z,\xi)\end{bmatrix}$.
Then, as outlined in the solution method in [37, Section 3], we can find stationary points of the bilevel program by solving the following system,
$\zeta\geq 0,\,\mathcal{G}(z,\zeta)\leq 0,\,\zeta^{T}\mathcal{G}(z,\zeta)=0,\,\mathcal{H}(z,\zeta;B_{i})=0.$
This system can be converted into a system of equalities via the Fischer-Burmeister function,
$\Phi(z,\zeta;B_{i})=\begin{pmatrix}\mathcal{H}(z,\zeta;B_{i})\\ \varphi_{\textup{FB}}\left(\zeta,-\mathcal{G}(z,\zeta)\right)\end{pmatrix}=0,$ (9)
where $\xi:=(z,\zeta)\in\mathbb{R}^{p}\times\mathbb{R}^{m\times p}\times\mathbb{R}^{2b+1}$ is a block variable and $\varphi_{\textup{FB}}$ is the Fischer-Burmeister, defined as
$\varphi_{\textup{FB}}(a,b)=\sqrt{a^{2}+b^{2}}-a-b.$
This system can then be solved using the Levenberg-Marquardt algorithm given in [37, Algorithm 1], where $\Psi:(\mathbb{R}^{p}\times\mathbb{R}^{mp})\times\mathbb{R}^{2b+1}$ is defined as,
[…]
The resulting weights of this solution can be used to construct a prediction function that has accounted for adversarial influence during the training process. With the inclusion of the constraints in (6), we can be confident that the adversary’s data retains the intended nature and realism after the transformation.
3 Numerical experiments
We assess the resilience of the bilevel model to strategic adversarial modifications. To do so, we design experiments on two datasets, the first of which, named Wine Quality records 11 features of 4898 bottles of red wine such as acidity and alcohol content. A corresponding quality score between 0100-10 is measured and recorded for each bottle. For this dataset, we consider scenarios where wine producers might lie when reporting the features of the wine or perhaps bribe quality assessors to alter the results in order to appear to be producing wine of a higher quality than they actually are. The second dataset, named Real Estate, contains the values of 6 features of 414 houses such as their age and distance to the nearest train station along with the corresponding sale price. We consider scenarios where sellers might lie about features of the house in order to be able to charge a higher price than deserved.
Further to this, we use the numerical experiments in this section as an opportunity to explore a particularly useful aspect of the constrained model which was not exploitable in [37]. Specifically, we demonstrate how, unlike existing pessimistic bilevel approaches to adversarial training, our models allow us to investigate which features of the adversary’s data were changed and hence give us insight into the most vulnerable aspects of the prediction models. In the current proposal for a pessimistic bilevel approach to evasion attacks, [5], pass the adversary’s data through a feature map. This plays a vital role in the proposed solution process, as the resulting feature space can be assumed to be unrestricted, which is convex. This, coupled with a strictly convex objective function, grantees a unique solution to the adversary’s problem in the lower-level, allowing the model to be reorganised as a pessimistic bilevel optimisation problem, which is easier to solve. However, since our model makes no assumptions on the convexity of the adversary’s problem or the uniqueness of its solution, we can leave the adversary’s data in its original feature space. Therefore, we can observe the adversary’s solution and infer the extent to which features must be changed in order to trick the learner. This analysis could not be performed in the experiments in [37], as the text-based data needed to be embedded before analysis. This embedding is not bijective and, consequently, the value of the adversary’s data at the solution provided no useful insight into the adversary’s strategy. However, in this chapter, we take the opportunity to test the bilevel model on some datasets that require no feature manipulation and hence allow us to investigate the adversary’s movements.
We divide the data into a training, and test set by the ration of $80\%$ and $20\%$ respectively and simulate adversarial attacks to inject the into the test set. Let $X^{\text{test}}\in\mathbb{R}^{T\times p}$ be the test set containing $T$ samples. Before evaluation, we simulate adversarial influence on the test set by transforming a portion of the instances towards target values $Z^{\text{test}}=Y^{\text{test}}+\Delta$, defined as a perturbation of the ground truth test labels $Y^{\text{test}}$, where $\Delta\in\mathbb{R}$ is a perturbation. Let $t<=T$ be the portion of data modified by an adversary and let $I\subset\{1,\dots,T\}$ be the corresponding indexes of the test data which are manipulated. We simulate a range of adversarial abilities. We generate adversarial test data by modifying $X^{\text{test}}_{I}$ towards the adversary’s target labels. To simulate a range of abilities, we randomly generate each adversary their own similarity threshold from the range $(0.8,1)$. Let $\delta^{\text{test}}\in(-1,1)^{t}$ be the set of similarity thresholds where $\delta_{i}\sim U(0.8,1)\;\forall\;i\in I$ and $U$ denotes the uniform distribution. The instances of adversarial test data and then found by solving the problem
$X^{\text{test}}_{i}:=\operatorname*{argmin}_{x\in\mathbb{R}^{q}}\left\{\ell(\sigma(w,x),Z^{\text{test}}_{i})\,\mid\,g(x,x^{0};\delta^{\text{test}}_{i})\leq 0\right\},\quad i\in I,$
where $x^{0}\in\mathbb{R}^{q}$ is the initial values of $x$. We set $T$ to be $10\%$ of the size of the test set. All features and labels are normalised to the range $(0,1)$ and we set the adversarial perturbation to be $\Delta=2\text{std}(Y^{\text{test}})$ where $\text{std}(Y^{\text{test}})$ is the standard deviation of the test labels. We compare our model to that of a typically trained linear regression predictor, named LinReg, for various values of $m$ and plot in Figure 1. We can see for both datasets a similar trend of an initially high mean square error (MSE) for $m=1$, before a drastic decrease to an optimal value of $m=2$. Following this, we see a gradual increase in MSE as $m$ increases, before plateauing around an MSE roughly equal to that of the traditional predictor.
Figure 1: Comparison of the pessimistic bilevel model with existing solvers
We now investigate the movement of each feature by the adversary during the training process to gain insight into which features are most vulnerable. For each feature (column), we calculate the average absolute distance between the adversary’s solution $X^{*}$ and the original position $X^{0}$,
$\frac{1}{m}\sum_{j=1}^{m}\left|X^{*}_{ij}-X^{0}_{ij}\right|,\quad i=1,\dots,m.$
Under this measure, a larger distance implies that the adversary needed to transform the data by a larger extent in order to sufficiently fool the leaner’s prediciton model. We plot the average movements for each feature and for each dataset in Figure 2. For the Wine quality, we see the most movement in alcohol followed closely by volatile acidity and residual sugar. These results give the insight into how much an adversary must modify each feature in order to fool the learner. It is clear, for example, that the alcohol content must be modified considerably more than the citric acid to generate the most effective wine. On the other hand, the citric acid content required will little change to achieve a value sufficient to fool the learner’s predictor. The citric acid, therefore, might be considered a particularly vulnerable feature that could easily be exploited and perhaps should be treated with more scrutiny. In the real estate data, we see the distance to the neaest MRT station as the most resilient feature, requiring relatively substantial change to effect the learner’s predictor. Meanwhile, the transaction data and longitude appear to be the easiest to exploit.
Figure 2: Feature modifications made by the adversary to trick the learner
It is clear from the experimental results that the constrained pessimistic bilevel model provides an effective method of accounting for adversarial manipulation during the training process of regression scenarios. The lower-level successfully anticipates the movements of an adversary, which leads to improved performance by the prediction model trained in the upper-level. We see a significant change in performance as we increase the number of instances available to be manipulated by the adversary. In particular, we note that a balance needs to be struck between allowing the adversary enough instances that they have sufficient influence over then training process, while not too many that the model overestimates the influence of the adversary. Finally, we demonstrated how, unlike the work in [37], we could investigate the values of the adversary’s data at the solution. This allowed us to identify which features were most vulnerable to exploitation.
4 Conclusion
Throughout this paper, we demonstrated how the pessimistic bilevel model proposed by [37] can be adapted to adversarial regression scenarios. In particular, by anticipating adversarial movement in the lower-level, we can train prediction functions that are resilient to adversaries who attempt to influence particular outcomes from a prediction model by strategically transforming their data. For example, in the context of property valuation, an adversary might purposefully provide falsified information to be achieve higher valuations and hence charge a higher selling price. While there exist pessimistic bilevel approaches to adversarial evasion scenarios, these models exploit strong assumptions about the convexity of the adversary’s problem and the uniqueness of their solution to reformulate the program into its optimistic variant, which, while easier to solve, oversimplifies the capabilities of the adversary. Since our model makes no such assumptions, we retain the pessimistic aspect of the bilevel model, allowing us to accurately capture the antagonistic nature of these scenarios.
To assess the again simulated adversarial influence on a number of applications, including insurance appraisal, wine quality appraisal and real estate pricing. Throughout numerical experiments, we demonstrated the ability of this model to train resilient predictors that provided improved performance over traditionally-trained predictors. Further to this, we again investigated varying the number of samples available to the adversary and the similarity threshold and observed a clear pattern in the form of a trade-off between allowing the adversary enough freedom and influence to generate sufficient movement that impacts the training process, while not too much such that the adversary provides too much influence over the training process, leading to overly pessimistic predictors.
Appendix A Derivatives for the leader and follower
Let $x\in\mathbb{R}^{q}$ be a sample of data with corresponding label $y\in\mathbb{R}$ abd let $w\in\mathbb{R}^{q}$ be the weights of the learner’s prediction function, $\sigma:\mathbb{R}\times\mathbb{R}\rightarrow\mathbb{R}$ as,
$\sigma(w,x)=w^{T}x.$
We define the upper-level (learner’s) loss function $\mathcal{L}:\mathbb{R}\times\mathbb{R}\rightarrow\mathbb{R}$ as,
$\mathcal{L}(\sigma(w,x),y):=\|\sigma(w,x)-y\|_{2}^{2}.$
Let $z\in\mathbb{R}$ be the adversary’s target label. We define the lower-level (adversary’s) loss function $\ell:\mathbb{R}\times\mathbb{R}\rightarrow\mathbb{R}$ as,
$\ell(\sigma(w,x),y):=\|\sigma(w^{T}x)-z\|_{2}^{2}.$
Let $D\in\mathbb{R}^{mq}$ be the static data with corresponding labels $\gamma\in\mathbb{R}^{n}$ and let $X\in\mathbb{R}^{mq}$ be the adversary’s data with corresponding labels $Y\in\mathbb{R}^{m}$. The upper-level objective $F:\mathbb{R}^{q}\times\mathbb{R}^{mq}\rightarrow\mathbb{R}$ is defined as,
$F(w,X):=\frac{1}{n}\sum_{i=1}^{n}\mathcal{L}(\sigma(w,D_{i}),\gamma_{i})+\frac{1}{m}\sum_{i=1}^{m}\mathcal{L}(\sigma(w,X_{i}),Y_{i})+\frac{1}{\rho}\|w\|_{2}^{2}.$
The derivative of the upper-level (leader’s) loss with respect to the learner’s weights is given by,
$\frac{\partial F}{\partial w}(w,X)=\frac{2}{n}D^{T}(w^{T}D-\gamma)+\frac{2}{m}X^{T}(w^{T}X-Y)+\frac{2}{\rho}w.$
The second derivative of the upper-level loss with respect to the learner’s weights is given by
$\frac{\partial^{2}F}{\partial ww}(w,X)=\frac{2}{n}D^{T}D+\frac{2}{m}X^{T}X+\frac{2}{\rho}$
The derivative of the upper-level loss function with respect to the adversary’s data is given by
$\frac{\partial\mathcal{L}(w,X)}{\partial X_{ij}}=\frac{2}{m}w_{j}(w^{T}X_{i}-y_{i}),\quad i=1,\dots,m,\;j=1,\dots,q.$
The second derivative with respect to the adversary’s data is given by the following cases,
$\frac{\partial^{2}\mathcal{L}}{\partial X_{ij}X_{kl}}=\begin{cases}\frac{2}{m}w_{j}w_{l}&i=k\\ 0&\text{otherwise}\end{cases}.$
Let $X^{0}\in\mathbb{R}^{mq}$ be the start point of the adversary’s data. We define the constraints $g:\mathbb{R}^{mq}\rightarrow(-1,1)$ of the lower-level component-wise as
$g_{i}(X)=\delta-d(X_{i},X_{i}^{0}),\;i\in\{1,\dots,m\},$
where
$d(X_{i},X_{i}^{0})=\frac{X_{i}\cdot X_{i}^{0}}{\|X_{i}\|\|X_{i}^{0}\|},\;i\in\{1,\dots,m\},$
and $\delta\in\mathbb{R}$ is the similarity threshold. The derivative of the constraints with respect to the classifier weights is 0. The derivative with respect to the adversary’s data is given by the following cases,
$\frac{\partial g_{i}(X)}{\partial X_{jk}}=\begin{cases}\frac{X_{ik}^{0}}{\|X_{i}\|\cdot\|X_{i}^{0}\|}-d(X_{i},X_{i}^{0})\frac{X_{ik}}{\|X_{i}\|^{2}}&i=j\\ 0&i\neq j\end{cases}.$
The second derivative with respect to the adversary’s data is given by the cases,
$\frac{\partial g_{i}(X)}{X_{jk}X_{lc}}=\begin{cases}\frac{X_{ic}X_{ik}^{0}+X_{ik}X_{ic}^{0}}{\|X_{i}\|^{3}\|X_{i}^{0}\|}-\frac{3X_{ik}X_{ic}d(X_{i},X_{i}^{0})}{\|X_{i}\|^{4}}&i=j=l,k\neq c\\ \frac{2X_{ik}X_{ik}^{0}}{\|X_{i}\|^{3}\|X_{i}^{0}\|}-\frac{3X_{ik}^{2}d(X_{i},X_{i}^{0})}{\|X_{i}\|^{4}}+\frac{d(X_{i},X_{i}^{0})}{\|X_{i}\|^{2}}&i=j=l,k=c\\ 0&\text{otherwise}.\end{cases}$

### v2

@@
 Adversarial machine learning challenges the assumption that the underlying distribution remains consistent throughout the training and implementation of a prediction model.
-In particular, adversarial evasion considers scenarios where adversaries adapt their data to influence particular outcomes from established prediction models, such scenarios arise in applications such as spam email filtering, malware detection and fake-image generation, where security methods must be actively updated to keep up with the ever-improving generation of malicious data.
+In particular, adversarial evasion considers scenarios where adversaries adapt their data to influence particular outcomes from established prediction models; such scenarios arise in applications such as spam email filtering, malware detection and fake-image generation, where security methods must be actively updated to keep up with the ever-improving generation of malicious data.
 Game theoretic models have been shown to be effective at modelling these scenarios and hence training resilient predictors against such adversaries.
-Recent advancements in the use of pessimistic bilevel optimsiation which remove assumptions about the convexity and uniqueness of the adversary’s optimal strategy have proved to be particularly effective at mitigating threats to classifiers due to its ability to capture the antagonistic nature of the adversary.
-However, this formulation has not yet been adapted to regression scenarios.
-This article serves to propose a pessimistic bilevel optimisation program for regression scenarios which makes no assumptions on the convexity or uniqueness of the adversary’s solutions.
+In this paper, we introduce a pessimistic bilevel optimsiation model, based on Stackelberg leader-follower games, to counter adversarial evasion in regression analysis.
+Unlike in some existing literature, we do not make assumptions (such as lower-level convexity or uniqueness of optimal solutions) on the adversary’s optimal strategy to avoid hindering their capacity and hence capture their antagonistic nature, leading to more resilient prediction models.
+Further to this, through the introduction of lower-level constraints, we measure and restrict the adversary’s movement, which, unlike its unrestricted counterpart in the current literature, prevents drastic transformations that do not accurately represent reality.
+Keywords: Adversarial learning, game theory, bilevel optimisation, regression
 1 Introduction
@@
 The way the adversary influences the distribution can fall under numerous categories, see [16] for a helpful taxonomy that categorises these attacks.
-We focus on the specific case of exploratory attacks, which consider the scenarios where adversaries attempt to modify their data to evade detection by a classifier.
+We focus on the specific case of evasion attacks, which consider the scenarios where adversaries attempt to modify their data to influence particular outcomes from prediction models.
 Such attacks might occur in security scenarios such as malware detection [2] and network intrusion traffic [23].
@@
 By modelling the adversary’s behaviour and anticipating these attacks, we can train classifiers that are resilient to such changes in the distribution before they occur.
-Game theory provides an effective and popular technique to model adversarial scenarios.
+Game theory provides an effective and popular technique to model adversarial evasion scenarios.
 These games see one player, the learner, attempt to train a prediction model while another player, the adversary, modifies their data in an attempt to influence particular outcomes from the learner’s predictor.
-See, for example [10] for early work in modelling adversarial learning.
+See, for example [7] for early work in modelling adversarial machine learning.
 The precise structure of such a model then depends on a number of factors, such as whether the attack occurs either at training time [21], or implementation time [4].
 Further to this, while some games assume the players act simultaneously [3, 22], others allow for sequential play where either the adversary acts first [19, 20, 5, 18] or the learner acts first [16, 4].
-We focus our attention on attacks made at implementation time where the adversary seeks to influence an already established prediction model, commonly known as evasion attacks in the literature [21].
+We focus our attention on attacks made at implementation time, where the adversary seeks to influence an already established prediction model.
 Sequential evasion games naturally take on the form of a bilevel optimisation program.
@@
 When multiple optimal solutions are available to the adversary, the pessimistic model, unlike its optimistic variant, assumes no cooperation between the adversary and the learner.
-This formulation better captures the antagonistic nature of adversarial evasion scenarios [2, 5].
-In this article, we adapt the pessimistic bilevel program proposed by [37] to train resilient predictors in regression scenarios.
-We model scenarios where we might expect adversaries to modify their data in an attempt to influence a particular label from a prediction model.
-We use the motivating example of an adversary providing falsified information about a house in order to influence a higher valuation.
-We explore properties of the constrained bilevel model under new choices of objective functions, showing the possible existence of multiple optimal solutions and hence the benefit of a constrained pessimistic bilevel approach, which, to the best of our knowledge, has not yet been applied to regression analysis.
-Moreover, we take this opportunity to exploit the fact that our method does not require the use of feature maps in its solution method.
-Feature maps, such as PCA, as was used [5], are typically not bijective.
-Consequently, although a solution to the adversary can be found in the map-induced feature space, a corresponding solution in the original feature space cannot be identified.
-However, under our model, we can view the explicit values of the adversary’s data at their solution.
+This formulation better captures the antagonistic nature of adversarial evasion scenarios [1, 27, 4].
+A pessimistic bilevel formulation for evasion attacks was initially proposed in [4] to suit classification tasks, which was later extended in [1, 27] by relaxing assumptions on convexity and uniqueness of solutions, again for classification problems.
+In this article, we present a novel pessimistic bilevel program to train resilient predictors against evasion attacks in regression problems.
+We consider scenarios where we expect adversaries to modify their data in an attempt to influence a particular label from a prediction model.
+For example, consider a real estate surveyor who uses a prediction model to evaluate the selling price of houses.
+A homeowner might provide falsified information about their house, such as it’s size or distance from public services, in order to influence a higher valuation from the surveyor.
+Consider a simple example of two houses, $H_{1}$ and $H_{2}$.
+These houses are identical in every aspect except for their location.
+While $H_{1}$ is located just 11km from the nearest train station, $H_{2}$ is located further away at 44km.
+Consequently, while $H_{1}$ is evaluated at £500500k, $H_{2}$ would receive a considerably lower evaluation due to its inaccessibility to public transport.
+To maximise the selling price, the owner of $H_{2}$ might lie about the distance of their house as closer to the train station than it actually is.
+Decisions like this are simulated by the adversary in the lower-level of our bilevel program.
+Then, the upper-level trains the surveyor’s prediction model while considering these adversarial influences and hence trains a more resilient prediction model.
+Within our model, we make no assumptions about the convexity of the lower-level problem or the uniqueness of its solution, retaining the benefit provided by the pessimistic formulation.
+Further to this, we introduce constraints on the lower-level optimisation problem which restrict the extent to which the adversary can modify their data.
+Consider again our simple housing market of houses $H_{1}$ and $H_{2}$.
+We established that the owner of $H_{2}$ can lie about the distance of their house from the nearest train station in order to gain a higher evaluation form the surveyor.
+Without restrictions on their movement, the owner of $H_{2}$ would simply state that their house is next to the train station, since a smaller distance leads to a higher evaluation.
+However, it is incredibility easy to spot this lie due to the large discrepancy between the true distance and the distance stated by the owner.
+A smaller change in distance, on the other hand, could enable the owner to receive a higher evaluation price while remaining plausibly realistic enough that the surveyor does not notice.
+Clearly, an adversary which takes smaller movements better reflects the strategies played by adversaries in the real world.
+However, without restrictions on their movement, the adversary will construct data which, while optimising their objective function, is not plausible or possibly nonsensical.
+The constraints we introduce here ensure that the similarity between the modified data and its original value is above some pre-defined threshold.
+To the best of our knowledge, existing game-theoretic approaches to adversarial regression analysis, see, for example, [28, 29, 30], do not propose a pessimistic bilevel program with lower-level constraints on the adversary’s movements that makes no assumptions on convexity or uniqueness of solutions.
+Moreover, where existing pessimistic approaches rely on feature maps, such as principal component analysis (PCA), or contextual embeddings to represent the adversary’s data, the adversary is allowed to manipulate their data in its original feature space.
+Consequently, we can view the explicit values of the adversary’s data at their solution.
 These values can give us insight into the strategies played by the adversary and allows us to identify the extent to which each feature needs to be modified in order to adequately influence the prediction model.
 From this, we can identify which features are most vulnerable.
-For example, if a certain feature required only a minor perturbation, then we might consider this feature to be particularly vulnerable to adversarial influence.
-On the other hand, if a feature required large transformations to trick the learner, then we might consider this feature to be resilient to attacks.
+For example, if a certain feature required only a minor perturbation, then we might consider it to be particularly vulnerable to adversarial influence.
+On the other hand, if a feature required large transformations to trick the learner, then we might consider it to be considerably more resilient to attacks.
 The contributions of this article can be summarised as follows.
-We adapt the bilevel model proposed in [37] to accommodate regression analysis and present a novel pessimistic bilevel optimisation program to train prediction functions that are resilient to strategic changes in the data of regression scenarios.
-This bilevel program makes no assumptions on the convexity of the lower-level problem, or the uniqueness of its solution, which, to the best of our knowledge, has not previously been proposed for adversarial regression analysis.
+We present a novel pessimistic bilevel optimisation program to train regression prediction functions that are resilient to strategic evasion attempts.
+This bilevel program, unlike existing approaches to adversarial regression, restricts the adversary’s movement through lower-level constraints, while also making no assumptions on the convexity of the lower-level problem, or the uniqueness of its solution.
 We design experiments to assess the performance of our model and showcase its ability to outperform existing methods.
-Finally, we demonstrate a key benefit of our model over existing approaches.
-Where existing approaches required the use of kernel transformations to ensure convexity, we are able to keep the lower-level variable (the adversary’s data) in its original feature space, allowing us to investigate the value of the data at the solution.
-From this we can infer which features require the greatest transformation in order to effectively influence the learner’s prediction function.
-The remainder of this paper is outlined as follows; we begin in Section 2 by presenting the constrained pessimistic bilevel model to train resilient prediction functions in regression analysis before outlining the solution method.
+Finally, we demonstrate a key benefit of our model over existing pessimistic approaches.
+Where existing approaches transform the adversary’s data before analysis, removing the ability to analyse the adversary’s data at their solution, we keep it in its original feature space.
+This grants us the ability to infer which features require the greatest transformation in order to effectively influence the learner’s prediction function.
+The remainder of this article is outlined as follows; we begin in Section 2 by presenting the pessimistic bilevel model with lower-level constraints to train resilient prediction functions in regression analysis before outlining the solution method.
 Following this, we assess the performance of the bilevel model on two datasets in Section 3 before concluding our findings in Section 4.
 2 Training resilient prediction functions
-In this section we present an adaptation of the constrained pessimistic bilevel program proposed in [37] to model adversarial regression.
+In this section we present the pessimistic bilevel program with lower-level constraints to model adversarial regression.
 In particular, we demonstrate how the process of adversarial training can be used to train resilient predictors under the scenario of adversaries attempting to strategically modify their data in an attempt to be assigned incorrect labels from established prediction models.
-Such scenarios might arise, for example, in property valuation.
-An adversary might lie and provide falsified information in an attempt to influence a higher valuation.
+Such scenarios might arise, for example, in property valuation, as outlined in the introduction.
 Through suitable choices of objective functions, we construct a constrained pessimistic bilevel program that sees a learner, in the upper-level, train a prediction function while considering how an adversary, in the lower-level, might alter their data towards a target outcome from the prediction model.
-The adversary is assigned some existing data as a starting point.
-The constraints are then used to measure the similarity between the adversary’s solution and starting point.
-As demonstrated in [37], these constraints ensure that the adversary’s data remain plausibly realistic and prevent instances of data losing their intended meaning.
+The adversary is assigned some existing data as a starting point and the distance that the adversary diverges from this point is measured and restricted by lower-level constraints.
+We demonstrate, through a motivating example, how these constraints ensure that the adversary’s data remain plausibly realistic and prevent instances of data losing their intended meaning.
 In this way, the bilevel model more realistically simulates adversarial transformations, leading to improved performance.
-To the best of our knowledge, a constrained pessimistic bilevel program with no assumptions on convexity or uniqueness of lower-level solution has previously been proposed for adversarial evasion scenarios in regression analysis.
+To the best of our knowledge, a pessimistic bilevel program with constrained adversarial movement and with no assumptions on convexity or uniqueness of lower-level solution has not previously been proposed for adversarial evasion scenarios in regression analysis.
 Let $x\in\mathbb{R}^{q}$ be a sample (row) of data containing $p\in\mathbb{N}$ features with corresponding label $y\in\mathbb{R}$.
-For some weights $w\in\mathbb{R}^{q}$, we define the learner’s prediction function, $\sigma:\mathbb{R}\times\mathbb{R}\rightarrow\mathbb{R}$, as the linear weighted sum of the features,
-$\sigma(w,x):=w_{0}+w^{T}x.$ (1)
-This prediction function is a common choice, for example, in linear regression.
+For some weights $w\in\mathbb{R}^{q}$, the learner’s prediction function, $\sigma:\mathbb{R}\times\mathbb{R}\rightarrow\mathbb{R}$, as the linear weighted sum of the features:
+$\sigma(w,x):=w^{T}x.$ (1)
+This is a common prediction function in linear regression.
 We then construct a loss function, $\mathcal{L}:\mathbb{R}\times\mathbb{R}\rightarrow\mathbb{R}$, to train the prediction function and identify appropriate weights.
 As in linear regression, we can achieve this by minimising the squared error between the learner’s prediction of $x$ and the label, $y$,
-$\mathcal{L}(\sigma(w,x),y):=(\sigma(w,x)-y)^{2}=(w_{0}+w^{T}x-y)^{2}.$
+$\mathcal{L}(\sigma(w,x),y):=(\sigma(w,x)-y)^{2}=(w^{T}x-y)^{2}.$ (2)
 The adversary, in the lower-level, transforms some data with the aim of having the learner’s prediction function, $\sigma$ output a desired target label.
@@
 If, for example, the actual valuation price is $y\in\mathbb{R}$, the adversary might set their target, $z\in\mathbb{R}$, as $z=y+\nu$, where $\nu>0$.
-The lower-level (adversary’s) loss function is then defined as the squared distance between the learner’s prediction and the target label,
+The lower-level (adversary’s) loss function is then defined as the squared distance between the learner’s prediction and the target label:
 $\ell(\sigma(w,x),z):=(\sigma(w,x)-z)^{2}=(w^{T}x-z)^{2}.$ (3)
 With the loss functions established, we now work them into objective functions that accommodate a full dataset.
-We divide the training data into two sets, a static set as would typically be used in the training of a machine learning predictor, and then second set of data which can be manipulated by the adversary.
-Let $D\in\mathbb{R}^{nq},D=(D_{1},\dots,D_{n})^{T}$ be the static set of $n\in\mathbb{N}$ instances of data where each $D_{i}\in\mathbb{R}^{q},\;i=1,\dots,n$, is a row vector containing the values of $q\in\mathbb{N}$ features, and let $\gamma\in\{0,1\}^{n}$ be the corresponding corresponding collection of labels.
-Let $X\in\mathbb{R}^{mq}$ be the set of $m\in\mathbb{N}$ instances which can be manipulated by the adversary, defined as,
+We divide the training data into two sets, a static set as would typically be used in the training of a machine learning predictor, and then a second set of data which can be manipulated by the adversary.
+Let $D\in\mathbb{R}^{n\times q}$ be the static set of $n\in\mathbb{N}$ instances of data where each $D_{i}\in\mathbb{R}^{q},\;i=1,\dots,n$, is a row vector containing the values of $q\in\mathbb{N}$ features, and let $\gamma\in\mathbb{R}^{n}$ be the corresponding corresponding collection of labels.
+Let $X\in\mathbb{R}^{m\times q}$ be the set of $m\in\mathbb{N}$ instances of the same $q$ features which can be manipulated by the adversary with corresponding labels $Y\in\{0,1\}^{m}$.
+For convenience, we collapse the adversary’s data into a single column vector in order to reduce its dimensionality.
+Therefore, the adversary’s data, $X\in\mathbb{R}^{mq}$ is refined as
 $X:=\begin{pmatrix}X_{1}^{T}\\ \vdots\\ X_{m}^{T}\end{pmatrix}$,
-where each $X_{i},\;i=1,\dots,m$, is a row vector of features.
-Let $Y\in\{0,1\}^{m}$ be the corresponding labels of the adversary’s data.
+where each $X_{i},\;i=1,\dots,m$, is a row vector of features. For consistency, we also redefine the static data in the same way, $D\in\mathbb{R}^{nq},D:=(D_{1},\dots,D_{n})^{T}$.
 The upper level objective of the bilevel program sees the learner minimising their loss over both sets of data.
-Note that the learner considers the true labels of the adversary’s data, where $F:\mathbb{R}^{p}\times\mathbb{R}^{mp}\rightarrow\mathbb{R}$ is the upper-level objective,
-$F(w,X):=\frac{1}{n}\sum_{i=1}^{n}\mathcal{L}(\sigma(w,D_{i}),\gamma_{i})+\frac{1}{m}\sum_{i=1}^{m}\mathcal{L}(\sigma(w,X_{i}),Y_{i}).$
+Note that the learner considers the true labels of the adversary’s data, where $F:\mathbb{R}^{p}\times\mathbb{R}^{mp}\rightarrow\mathbb{R}$ is the upper-level objective defined by
+$F(w,X):=\frac{1}{n}\sum_{i=1}^{n}\mathcal{L}(\sigma(w,D_{i}),\gamma_{i})+\frac{1}{m}\sum_{i=1}^{m}\mathcal{L}(\sigma(w,X_{i}),Y_{i}).$ (4)
 Then, given a set of weights, $w$ of the learner’s prediction function, the adversary, in the lower-level seeks to modify their data such that the learner’s prediction moves towards their target labels, $Z$.
-Let $f:\mathbb{R}^{p}\times\mathbb{R}^{mp}\rightarrow\mathbb{R}$ be the adversary’s objective, which is defined as the sum of the adversary’s loss over their data,
+Let $f:\mathbb{R}^{p}\times\mathbb{R}^{mp}\rightarrow\mathbb{R}$ be the adversary’s objective, which is defined as the sum of the adversary’s loss over their data:
 $f(w,X):=\frac{1}{m}\sum_{i=1}^{m}\ell(\sigma(w,X_{i}),Z_{i}).$ (5)
-Finally, we define the constraints $g:\mathbb{R}^{mq}\rightarrow\mathbb{R}^{m}$ as
-$g(X):=\begin{pmatrix}g_{1}(X_{1})\\ \vdots\\ g_{m}(X_{m})\end{pmatrix}$, (4)
-where each constraint function, $g_{i}:\mathbb{R}^{q}\rightarrow\mathbb{R},\;i=1,\dots,m$, is defined as
-$g(X):=\delta-d(X_{i},X^{0}_{i}),$ (5)
-where $d:\mathbb{R}^{q}\times\mathbb{R}^{q}\rightarrow\mathbb{R}$ is some similarity function.
-These constraints serve to measure and restrict the cosine similarity between the adversary’s solution and initial value to be below some threshold $\delta\in\mathbb{R}$.
-As outlined in [37], these constraints prevent the adversary’s data becoming nonsensical and too far removed from its intended message.
-This continues to be useful under the regression scenarios as we can reduce the likelihood of the adversary producing unrealistic data.
-For example, in reality, an adversary cannot claim that their house was built thousands of years ago when in reality it was built a mere 5 years ago.
-This would be immediately suspicious to anybody who checks the adversary’s data.
-However, smaller transformations, such as within the range of $\pm 2$ years would be far more believable.
-Similarly to the work in [37], we use the cosine similarity to measure the extent to which the adversary has modified their data.
-Let $X^{0}\in\mathbb{R}^{mq}$ be the initial value of the adversary’s data, then the constraints are then defined as follows,
-$g_{i}\left(X\right)=\delta-d(X_{i},X^{0}_{i})=\delta-\frac{X_{i}\cdot X_{i}^{0}}{\|X_{i}\|\|X_{i}^{0}\|},\;i\in\{1,\dots,m\}.$ (6)
-Then, setting the initial value of the adversary’s data to some existing data in the training set, we can be confident that under these constraints, the adversary produces data that are similar in nature to some real example.
-We now formulate the interactions between the learner and adversary as a bilevel optimisation problem.
-To begin, given a set of weights, $w$, of the learner’s prediction function, let $S(w)$ be the set of optimal data for the adversary,
-$S(w):=\operatorname*{argmin}_{X\in\mathbb{R}^{mq}}\left\{f(w,X)\mid g(X_{i})\leq 0\right\},$ (7)
-This set contains all instances of data that will trick the learner into assigning the adversary their desired labels $Z$.
-For example, the falsified falsified report that gives a property a better valuation than it actually deserves.
+As we have highlighted through a motivating real estate example in Section 1, an adversary who takes smaller movements better reflects the strategies played by adversaries in the real world.
+However, the unconstrained objective functions do not take this under consideration.
+The optimal adversary under the objective function in (5) will overestimate the abilities of the adversary, leading unrealistic data transformations.
+In order to construct a realistic adversary within the bilevel program, we introduce a set of lower-level constraints on the adversary’s optimization problem which measure the similarity between the true value of the adversary’s data and the value seen by the learner.
+Let $X^{0}$ be the true (original) value of the adversary’s data, and let $g:\mathbb{R}^{mq}\rightarrow\mathbb{R}^{m}$ be the vector of constraint functions
+$g(X):=\begin{pmatrix}g_{1}(X)\\ \vdots\\ g_{m}(X)\end{pmatrix}$. (6)
+We define each constraint function component $g_{i}:\mathbb{R}^{q}\rightarrow\mathbb{R}$, for each instance of the adversary’s data $i=1,\ldots,m$, by
+$g_{i}(X):=\delta-d(X_{i},X^{0}_{i}),$ (7)
+where $d:\mathbb{R}^{q}\times\mathbb{R}^{q}\rightarrow\mathbb{R}$ is some similarity function and $\delta$ is the minimum required similarity threshold.
+The constraint $g(X)\leq 0$ then restrict the adversary to produce data whose similarity is greater than $\delta$.
+The work in [27] demonstrated the effectiveness of the cosine similarity to measure the extent to which the adversary has modified their data and so we use the same here,
+$d(X_{i},X^{0}_{i}):=\frac{X_{i}\cdot X_{i}^{0}}{\|X_{i}\|\|X_{i}^{0}\|},\;i=\{1,\dots,m\}.$ (8)
+Combining the lower-level objective 5 with the constraints in 8 gives the adversary’s optimisation problem, the solution to which provides data that trick the learner into giving the best possible label assignment while remaining plausibly realistic.
+Given a set of weights of the learner’s prediction function, $w$, we defined $S(w)$ as this set of optimal data, given by
+$S(w):=\operatorname*{argmin}_{X\in\mathbb{R}^{mq}}\left\{f(w,X)\mid g(X)\leq 0\right\}.$ (9)
 Note that the adversary’s objective function under the linear regression loss in (3) and prediction function (1) possess multiple optimal solutions.
-This can be directly shown using in [37, Proposition 1] since the prediction function, $\sigma$, can be expressed as a function of the linear combination $w_{0}+w^{T}x$ for some $x\in\mathbb{R}^{q}$.
-Simply let $\Omega(w_{0}+w^{T}x)=w_{0}+w^{T}x$ and the proof follows.
-Note also that its has already demonstrated in [37, Proposition 2] that the feasible region defined by these constraints is non-convex.
+This can be directly shown using [27, Proposition 1] since the prediction function, $\sigma(w,x)=w^{T}x$, is a linear combination of $w$ and $x$.
+Note also that it has already been demonstrated in [27, Proposition 2] that the feasible region defined by the constraints in (8) is non-convex.
 Therefore, in the case of multiple optimal solutions to the adversary’s problem, given by the set $S(w)$, we make the pessimistic assumption that the adversary will not act cooperatively, namely, they will choose the solution which maximises the learner’s objective, while the learner seeks to minimise it.
@@
 $\min_{w\in\mathbb{R}^{p}}\max_{X\in S(w)}F(w,X).$ (10)
-The non-convexity and existence of multiple optimal solutions to the lower-level problem motives the use of the solution method proposed in [37].
-To begin, we collect the upper and lower variables, $w$ and $X$, into the block variable $z$ as below
-$z:=\begin{bmatrix}w\\ X\end{bmatrix}\in\mathbb{R}^{q+mq},\qquad$
-and define the block variable $\zeta\in\mathbb{R}^{2b+1}$ as,
-$\xi:=\begin{bmatrix}\beta\\ \hat{\beta}\\ \lambda\end{bmatrix}\in\mathbb{R}^{2b+1},$
-where $\beta,\hat{\beta}\in\mathbb{R}^{b}$ and $\lambda\in\mathbb{R}$ are lagrange multiples with associated Lagrangian-type functions $L^{\textup{p}}_{w},L^{\textup{p}}_{X},\ell^{\textup{p}}\colon(\mathbb{R}^{q}\times\mathbb{R}^{mq})\times\mathbb{R}^{2b+1}\to\mathbb{R}$ as stated below:
-$L^{\textup{p}}_{w}(z,\xi):=-F(w,X)+(\beta-\lambda\hat{\beta})^{\top}g(X)$
-$L^{\textup{p}}_{X}(z,\xi):=-F(w,X)+\lambda f(w,X)+\beta^{\top}g(X)$
+A solution to the problem described by (9) - (10) will comprise the weights of a prediction that has accounted for adversarial manipulation during its training process.
+Consequently, the prediction function will be more resilient to evasion attacks made after implementation.
+We say that a set of weights $\bar{w}$ is a local optimal solution for problem (9) - (10) if there is a neighbourhood $W$ of the point such that
+$\varphi_{p}(\bar{w})\leq\varphi_{p}(w)\;\,\mbox{ for all }\;w\in W,$ (11)
+where the $\varphi_{p}$ denotes the following two-level value function (concept first introduced and studied in [10]):
+$\varphi_{\textup{p}}(w):=\max\{F(w,X)\,|\,X\in S(w)\}.$
+[…]
+$\nabla_{w}F(w,X)$ $=0,$ (12a)
+$\nabla_{X}F(w,X)-\lambda\nabla_{X}f(w,X)-\nabla g(X)^{\top}\beta$ $=0,$ (12b)
+$\nabla_{X}f(w,X)+\nabla g(X)^{\top}\hat{\beta}$ $=0,$ (12c)
+$\hat{\beta}\geq 0,\quad g(X)\leq 0,\quad\hat{\beta}^{\top}g(X)$ $=0,$ (12d)
+$\lambda\geq 0,\;\,\beta\geq 0,\quad g(X)\leq 0,\quad\beta^{\top}g(X)$ $=0,$ (12e)
+for some data $X$ belonging to the adversary. We can solve this system by first transforming it into a system of equations. To simplify the notation, we introduce the block variables
+$z:=\begin{bmatrix}w\\ X\end{bmatrix}\in\mathbb{R}^{q+mq}\;\mbox{ and }\;\xi:=\begin{bmatrix}\beta\\ \hat{\beta}\\ \lambda\end{bmatrix}\in\mathbb{R}^{2b+1},$
+as well as the block functions
+$\mathcal{G}(z):=\begin{bmatrix}g(X)\\ g(X)\\ 0\end{bmatrix}\;\mbox{ and }\;\mathcal{H}(z,\xi):=\begin{bmatrix}\nabla_{w}F(w,X)\\ \nabla_{X}L^{\textup{p}}_{X}(z,\xi)\\ \nabla_{X}\ell^{\textup{p}}(z,\xi)\end{bmatrix}$,
+where $L^{\textup{p}}_{X},\,\ell^{\textup{p}}\colon(\mathbb{R}^{q}\times\mathbb{R}^{mq})\times\mathbb{R}^{2b+1}\to\mathbb{R}$ are upper and lower-level Lagrangian type functions:
+$L^{\textup{p}}_{X}(z,\xi):=F(w,X)-\lambda f(w,X)-\beta^{\top}g(X),$
 $\ell^{\textup{p}}(z,\xi):=f(w,X)+\hat{\beta}^{\top}g(X).$
-Now, we can define the block functions, $\mathcal{G}$ and $\mathcal{H}$ as below,
-$\mathcal{G}(z):=\begin{bmatrix}g(X)\\ g(X)\\ 0\end{bmatrix},\qquad\mathcal{H}(z,\xi;B_{i}):=\begin{bmatrix}\nabla_{w}L^{\textup{p}}_{w}(z,\xi)\\ \nabla_{X}L^{\textup{p}}_{X}(z,\xi)\\ \nabla_{X}\ell^{\textup{p}}(z,\xi)\end{bmatrix}$.
-Then, as outlined in the solution method in [37, Section 3], we can find stationary points of the bilevel program by solving the following system,
-$\zeta\geq 0,\,\mathcal{G}(z,\zeta)\leq 0,\,\zeta^{T}\mathcal{G}(z,\zeta)=0,\,\mathcal{H}(z,\zeta;B_{i})=0.$
-This system can be converted into a system of equalities via the Fischer-Burmeister function,
-$\Phi(z,\zeta;B_{i})=\begin{pmatrix}\mathcal{H}(z,\zeta;B_{i})\\ \varphi_{\textup{FB}}\left(\zeta,-\mathcal{G}(z,\zeta)\right)\end{pmatrix}=0,$ (9)
-where $\xi:=(z,\zeta)\in\mathbb{R}^{p}\times\mathbb{R}^{m\times p}\times\mathbb{R}^{2b+1}$ is a block variable and $\varphi_{\textup{FB}}$ is the Fischer-Burmeister, defined as
-$\varphi_{\textup{FB}}(a,b)=\sqrt{a^{2}+b^{2}}-a-b.$
-This system can then be solved using the Levenberg-Marquardt algorithm given in [37, Algorithm 1], where $\Psi:(\mathbb{R}^{p}\times\mathbb{R}^{mp})\times\mathbb{R}^{2b+1}$ is defined as,
-[…]
-The resulting weights of this solution can be used to construct a prediction function that has accounted for adversarial influence during the training process.
-With the inclusion of the constraints in (6), we can be confident that the adversary’s data retains the intended nature and realism after the transformation.
+Based on this notation, the system described by (12a)–(12e) can then be restated as
+$\zeta\geq 0,\;\;\mathcal{G}(z)\leq 0,\;\;\zeta^{T}\mathcal{G}(z)=0,\;\;\mathcal{H}(z,\xi)=0,$ (13)
+which can equivalently be written as the following system of equations
+$\left\{\begin{array}[]{l}\mathcal{H}(z,\zeta)=0,\\ \vartheta_{\textup{FB}}\left(\zeta_{i},-\mathcal{G}_{i}(z,\zeta)\right)=0,\;i=1,\ldots,2m.\end{array}\right$. (14)
+Here, $\vartheta_{\textup{FB}}$ corresponds to the so-called Fischer-Burmeister function [13], which is defined by
+$\vartheta_{\text{FB}}(a,b):=\sqrt{a^{2}+b^{2}}-(a+b)\;\mbox{ for }\;(a,b)\in\mathbb{R}^{2}.$
+Thanks to this function, the second equation in (13) is equivalent to the complementarity conditions in (12d)–(12e), written in compact form in (13), while considering them in pairs.
+Clearly, the stationary conditions (12a)–(12e) of problem (9) - (10) have been written as a system of equations in (14).
+This system is overdetermined, with $m$ more equations than variables, making the Levenberg-Marquardt method, specifically, the global nonsmooth Levenberg–Marquardt method for mixed nonlinear complementarity systems developed in [17], a suitable choice to solve it.
+A version of this algorithm which has been appropriately modified for the pessimistic adversarial bilevel program by introducing additional stopping criteria, is given in [27, Algorithm 2].
+We apply this same algorithm here to solve system (14), using the derivative formula provided in Appendix A, and extract the solution $z^{*}=(w^{*},X^{*})$.
+This solution comprises the weights, $w^{*}$, of a prediction function that has accounted for adversarial influence during the training process as well as the corresponding transformed adversary’s data, $X^{*}$.
+Unlike the framework proposed in the previous work in [27], it is not required that the adversary’s data is embedded ahead of analysis.
+Consequently, we can use $X^{*}$ to observe the value of the adversary’s solution in its original feature space and investigate the nature of their transformation.
 3 Numerical experiments
@@
 To do so, we design experiments on two datasets, the first of which, named Wine Quality records 11 features of 4898 bottles of red wine such as acidity and alcohol content.
-A corresponding quality score between 0100-10 is measured and recorded for each bottle.
+A corresponding quality score between 1101-10 is measured and recorded for each bottle.
 For this dataset, we consider scenarios where wine producers might lie when reporting the features of the wine or perhaps bribe quality assessors to alter the results in order to appear to be producing wine of a higher quality than they actually are.
@@
 We consider scenarios where sellers might lie about features of the house in order to be able to charge a higher price than deserved.
-Further to this, we use the numerical experiments in this section as an opportunity to explore a particularly useful aspect of the constrained model which was not exploitable in [37].
+Further to this, we use the numerical experiments in this section to explore a particularly useful aspect of the constrained model which, due to the requirement of feature transformations, was not exploitable in previous works.
 Specifically, we demonstrate how, unlike existing pessimistic bilevel approaches to adversarial training, our models allow us to investigate which features of the adversary’s data were changed and hence give us insight into the most vulnerable aspects of the prediction models.
-In the current proposal for a pessimistic bilevel approach to evasion attacks, [5], pass the adversary’s data through a feature map.
-This plays a vital role in the proposed solution process, as the resulting feature space can be assumed to be unrestricted, which is convex.
-This, coupled with a strictly convex objective function, grantees a unique solution to the adversary’s problem in the lower-level, allowing the model to be reorganised as a pessimistic bilevel optimisation problem, which is easier to solve.
-However, since our model makes no assumptions on the convexity of the adversary’s problem or the uniqueness of its solution, we can leave the adversary’s data in its original feature space.
-Therefore, we can observe the adversary’s solution and infer the extent to which features must be changed in order to trick the learner.
-This analysis could not be performed in the experiments in [37], as the text-based data needed to be embedded before analysis.
-This embedding is not bijective and, consequently, the value of the adversary’s data at the solution provided no useful insight into the adversary’s strategy.
-However, in this chapter, we take the opportunity to test the bilevel model on some datasets that require no feature manipulation and hence allow us to investigate the adversary’s movements.
 We divide the data into a training, and test set by the ration of $80\%$ and $20\%$ respectively and simulate adversarial attacks to inject the into the test set.
@@
 The instances of adversarial test data and then found by solving the problem
-$X^{\text{test}}_{i}:=\operatorname*{argmin}_{x\in\mathbb{R}^{q}}\left\{\ell(\sigma(w,x),Z^{\text{test}}_{i})\,\mid\,g(x,x^{0};\delta^{\text{test}}_{i})\leq 0\right\},\quad i\in I,$
+$X^{\text{test}}_{i}:=\operatorname*{argmin}_{x\in\mathbb{R}^{q}}\left\{\ell\left(\sigma(w^{\text{init}},x\right),Z^{\text{test}}_{i})\,\left|\,\delta^{\text{test}}_{i}-\frac{x\cdot Z^{\text{test}}_{i}}{\|x\|\|Z^{\text{test}}_{i}\|}\leq 0\right.\right\},\;\forall i\in I.$
 where $x^{0}\in\mathbb{R}^{q}$ is the initial values of $x$.
@@
 All features and labels are normalised to the range $(0,1)$ and we set the adversarial perturbation to be $\Delta=2\text{std}(Y^{\text{test}})$ where $\text{std}(Y^{\text{test}})$ is the standard deviation of the test labels.
-We compare our model to that of a typically trained linear regression predictor, named LinReg, for various values of $m$ and plot in Figure 1.
-We can see for both datasets a similar trend of an initially high mean square error (MSE) for $m=1$, before a drastic decrease to an optimal value of $m=2$.
-Following this, we see a gradual increase in MSE as $m$ increases, before plateauing around an MSE roughly equal to that of the traditional predictor.
+We compare our model to that of a typically trained linear regression predictor, named LinReg.
+While the pessimistic model in [4] was intended for use on classification tasks, its prediction function and loss functions can be easily substituted for linear regression variants which still satisfy their assumptions about strict convexity.
+We name this B&S and use it as a comparison to the existing pessimistic approach to adversarial regression analysis.
+The mean square error for these models as well as our model for various values of the adversary’s sample size, $m$, and the similarity threshold, $\delta$, are plotted in Figure 1.
+We can see for both datasets a similar trend of an initially high mean square error (MSE) for low values of $m$, before a decrease to an optimal value.
+Following this, we observe an increase in MSE, before plateauing around an MSE roughly equal to that of the traditional predictor.
+Both experiments see an optimal MSE when $\delta=0.95$ and when $m$ is fairly low.
+For the wine dataset, the optimal value falls at $m=1$, while on the real estate dataset, we see the optimal value fall at $m=2$.
+Although, we note that the real estate dataset also sees similarly good performance for $m$ in the range of 132313-23.
+In general, we observe a general pattern in that the adversary should be granted enough freedom and influence over the training process to capture the nature of adversarial attacks, while not too much influence that the model becomes over-estimates the power of the adversary and hence suffers in performance.
 Figure 1: Comparison of the pessimistic bilevel model with existing solvers
@@
 $\frac{1}{m}\sum_{j=1}^{m}\left|X^{*}_{ij}-X^{0}_{ij}\right|,\quad i=1,\dots,m.$
-Under this measure, a larger distance implies that the adversary needed to transform the data by a larger extent in order to sufficiently fool the leaner’s prediciton model.
+Under this measure, a larger distance implies that the adversary needed to transform the data by a larger extent in order to sufficiently fool the leaner’s prediction model.
 We plot the average movements for each feature and for each dataset in Figure 2.
-For the Wine quality, we see the most movement in alcohol followed closely by volatile acidity and residual sugar.
+For the Wine quality, we see the most movement in residual sugar followed by PH and sulfur dioxide.
 These results give the insight into how much an adversary must modify each feature in order to fool the learner.
-It is clear, for example, that the alcohol content must be modified considerably more than the citric acid to generate the most effective wine.
-On the other hand, the citric acid content required will little change to achieve a value sufficient to fool the learner’s predictor.
-The citric acid, therefore, might be considered a particularly vulnerable feature that could easily be exploited and perhaps should be treated with more scrutiny.
-In the real estate data, we see the distance to the neaest MRT station as the most resilient feature, requiring relatively substantial change to effect the learner’s predictor.
+It is clear, for example, that the alcohol content must be modified considerably to generate the most effective wine.
+While chlorides, on the other hand, required little change to achieve a value sufficient to fool the learner’s predictor.
+Chlorides, therefore, might be considered a particularly vulnerable feature that could easily be exploited and perhaps should be treated with more scrutiny.
+In the real estate data, we see the distance to the nearest MRT station as the most resilient feature, requiring relatively substantial change to affect the learner’s predictor.
 Meanwhile, the transaction data and longitude appear to be the easiest to exploit.
 Figure 2: Feature modifications made by the adversary to trick the learner
-It is clear from the experimental results that the constrained pessimistic bilevel model provides an effective method of accounting for adversarial manipulation during the training process of regression scenarios.
+It is clear from the experimental results that the pessimistic bilevel model with constrained adversarial movement provides an effective method of accounting for adversarial manipulation during the training process of regression scenarios.
 The lower-level successfully anticipates the movements of an adversary, which leads to improved performance by the prediction model trained in the upper-level.
@@
 In particular, we note that a balance needs to be struck between allowing the adversary enough instances that they have sufficient influence over then training process, while not too many that the model overestimates the influence of the adversary.
-Finally, we demonstrated how, unlike the work in [37], we could investigate the values of the adversary’s data at the solution.
+Finally, we demonstrated how, unlike existing works, we could investigate the values of the adversary’s data at the solution.
 This allowed us to identify which features were most vulnerable to exploitation.
 4 Conclusion
-Throughout this paper, we demonstrated how the pessimistic bilevel model proposed by [37] can be adapted to adversarial regression scenarios.
+In this article, we proposed a pessimistic bilevel program to model adversarial evasion in regression scenarios.
 In particular, by anticipating adversarial movement in the lower-level, we can train prediction functions that are resilient to adversaries who attempt to influence particular outcomes from a prediction model by strategically transforming their data.
-For example, in the context of property valuation, an adversary might purposefully provide falsified information to be achieve higher valuations and hence charge a higher selling price.
+For example, in the context of property valuation, an adversary might purposefully provide falsified information to achieve higher valuations and hence charge a higher selling price.
 While there exist pessimistic bilevel approaches to adversarial evasion scenarios, these models exploit strong assumptions about the convexity of the adversary’s problem and the uniqueness of their solution to reformulate the program into its optimistic variant, which, while easier to solve, oversimplifies the capabilities of the adversary.
 Since our model makes no such assumptions, we retain the pessimistic aspect of the bilevel model, allowing us to accurately capture the antagonistic nature of these scenarios.
-To assess the again simulated adversarial influence on a number of applications, including insurance appraisal, wine quality appraisal and real estate pricing.
-Throughout numerical experiments, we demonstrated the ability of this model to train resilient predictors that provided improved performance over traditionally-trained predictors.
-Further to this, we again investigated varying the number of samples available to the adversary and the similarity threshold and observed a clear pattern in the form of a trade-off between allowing the adversary enough freedom and influence to generate sufficient movement that impacts the training process, while not too much such that the adversary provides too much influence over the training process, leading to overly pessimistic predictors.
+Furthermore, with the introduction of lower-level constraints, which restrict the adversary’s movement, preventing nonsensical transformations and ensuring the adversary’s data remains plausibly realistic, leading to more accurate prediction functions.
+We simulated adversarial influence to create test sets for wine quality appraisal and real estate pricing.
+Numerical experiments demonstrated the ability of the bilevel model to train resilient predictors that provided improved performance over existing methods.
+Further to this, we investigated varying the number of samples available to the adversary and observed a clear pattern in the form of a trade-off between allowing the adversary enough freedom and influence to generate sufficient movement that impacts the training process, while not too much such that the adversary provides too much influence over the training process, leading to overly pessimistic predictors.
+Additionally, we investigated a feature of the pessimistic bilevel model which previous works were not able to exploit.
+Specifically, since we do not transform the adversary’s data before analysis, we were able to observe the adversary’s solution in its original feature space and measure the extent to which the adversary modified each feature.
+From this we could identify which features were most vulnerable to attacks.
 Appendix A Derivatives for the leader and follower
-Let $x\in\mathbb{R}^{q}$ be a sample of data with corresponding label $y\in\mathbb{R}$ abd let $w\in\mathbb{R}^{q}$ be the weights of the learner’s prediction function, $\sigma:\mathbb{R}\times\mathbb{R}\rightarrow\mathbb{R}$ as,
-$\sigma(w,x)=w^{T}x.$
-We define the upper-level (learner’s) loss function $\mathcal{L}:\mathbb{R}\times\mathbb{R}\rightarrow\mathbb{R}$ as,
-$\mathcal{L}(\sigma(w,x),y):=\|\sigma(w,x)-y\|_{2}^{2}.$
-Let $z\in\mathbb{R}$ be the adversary’s target label. We define the lower-level (adversary’s) loss function $\ell:\mathbb{R}\times\mathbb{R}\rightarrow\mathbb{R}$ as,
-$\ell(\sigma(w,x),y):=\|\sigma(w^{T}x)-z\|_{2}^{2}.$
+Let $x\in\mathbb{R}^{q}$ be a sample of data with corresponding label $y\in\mathbb{R}$ and let $w\in\mathbb{R}^{q}$ be the weights of the learner’s prediction function, $\sigma:\mathbb{R}\times\mathbb{R}\rightarrow\mathbb{R}$ defined as in (1).
+We define the upper-level (learner’s) loss function $\mathcal{L}:\mathbb{R}\times\mathbb{R}\rightarrow\mathbb{R}$ as it is in (2).
+Let $z\in\mathbb{R}$ be the adversary’s target label.
 Let $D\in\mathbb{R}^{mq}$ be the static data with corresponding labels $\gamma\in\mathbb{R}^{n}$ and let $X\in\mathbb{R}^{mq}$ be the adversary’s data with corresponding labels $Y\in\mathbb{R}^{m}$.
-The upper-level objective $F:\mathbb{R}^{q}\times\mathbb{R}^{mq}\rightarrow\mathbb{R}$ is defined as,
-$F(w,X):=\frac{1}{n}\sum_{i=1}^{n}\mathcal{L}(\sigma(w,D_{i}),\gamma_{i})+\frac{1}{m}\sum_{i=1}^{m}\mathcal{L}(\sigma(w,X_{i}),Y_{i})+\frac{1}{\rho}\|w\|_{2}^{2}.$
+The upper-level objective $F:\mathbb{R}^{q}\times\mathbb{R}^{mq}\rightarrow\mathbb{R}$ is defined as it is in (4).
 The derivative of the upper-level (leader’s) loss with respect to the learner’s weights is given by,
-$\frac{\partial F}{\partial w}(w,X)=\frac{2}{n}D^{T}(w^{T}D-\gamma)+\frac{2}{m}X^{T}(w^{T}X-Y)+\frac{2}{\rho}w.$
+$\nabla_{w}F(w,X)=\frac{2}{n}D^{T}(w^{T}D-\gamma)+\frac{2}{m}X^{T}(w^{T}X-Y)+\frac{2}{\rho}w.$
 The second derivative of the upper-level loss with respect to the learner’s weights is given by
-$\frac{\partial^{2}F}{\partial ww}(w,X)=\frac{2}{n}D^{T}D+\frac{2}{m}X^{T}X+\frac{2}{\rho}$
-The derivative of the upper-level loss function with respect to the adversary’s data is given by
+$\nabla^{2}_{ww}F(w,X)=\frac{2}{n}D^{T}D+\frac{2}{m}X^{T}X+\frac{2}{\rho}$
+The derivative of the upper-level objective function with respect to the adversary’s data is given by
+$\frac{\partial F}{\partial X_{ij}}(w,X):=\frac{1}{m}\sum_{k=1}^{m}\frac{\partial\mathcal{L}}{\partial X_{ij}}(\sigma(w,X_{k}),Y_{k})=\frac{1}{m}\frac{\partial\mathcal{L}}{\partial X_{ij}}(\sigma(w,X_{i}),Y_{i}),$
+where
 $\frac{\partial\mathcal{L}(w,X)}{\partial X_{ij}}=\frac{2}{m}w_{j}(w^{T}X_{i}-y_{i}),\quad i=1,\dots,m,\;j=1,\dots,q.$
 The second derivative with respect to the adversary’s data is given by the following cases,
-$\frac{\partial^{2}\mathcal{L}}{\partial X_{ij}X_{kl}}=\begin{cases}\frac{2}{m}w_{j}w_{l}&i=k\\ 0&\text{otherwise}\end{cases}.$
-Let $X^{0}\in\mathbb{R}^{mq}$ be the start point of the adversary’s data. We define the constraints $g:\mathbb{R}^{mq}\rightarrow(-1,1)$ of the lower-level component-wise as
-$g_{i}(X)=\delta-d(X_{i},X_{i}^{0}),\;i\in\{1,\dots,m\},$
-where
-$d(X_{i},X_{i}^{0})=\frac{X_{i}\cdot X_{i}^{0}}{\|X_{i}\|\|X_{i}^{0}\|},\;i\in\{1,\dots,m\},$
-and $\delta\in\mathbb{R}$ is the similarity threshold.
+$\frac{\partial^{2}\mathcal{L}}{\partial X_{ij}\partial X_{kl}}=\begin{cases}\frac{2}{m}w_{j}w_{l}&i=k\\ 0&\text{otherwise}.\end{cases}$
+The derivative of the upper-level objective function with respect to the adversary’s data is given by
+$\frac{\partial f}{\partial X_{ij}}(w,X)=\frac{1}{m}\sum_{i=k}^{m}\frac{\partial\ell}{\partial X_{ij}}(\sigma(w,X_{k}),Y_{k})=\frac{1}{m}\frac{\partial\ell}{\partial X_{ij}}(\sigma(w,X_{k}),Y_{k}).$
+where the lower-level (adversary’s) loss function $\ell:\mathbb{R}\times\mathbb{R}\rightarrow\mathbb{R}$ is defined as it is in (3).
+Note that the adversary’s loss function can be expressed in terms of the leader’s loss function,
+$\ell(\sigma(w,x),z)=\mathcal{L}(\sigma(w,x),z),$
+for some weights $w\in\mathbb{R}^{q}$, data $x\in\mathbb{R}^{q}$ and target label $z\in\mathbb{R}$.
+Therefore, we can express the derivative of the lower-level objective function with respect to the adversary’s data as follows
+$\frac{\partial f}{\partial X_{ij}}(w,X)=\frac{1}{m}\frac{\partial\mathcal{L}}{\partial X_{ij}}(\sigma(w,X_{i}),Z_{i})\quad i=1,\dots,m,\;j=1,\dots,q.$
+The second derivative with respect to the adversary’s data is then given as
+$\frac{\partial^{2}f}{\partial X_{ij}\partial X_{kc}}(w,X)=\frac{1}{m}\frac{\partial^{2}\mathcal{L}}{\partial X_{ij}\partial X_{jk}}(\sigma(w,X_{i}),Z_{i})\quad i,k=1,\dots,m,\;j,c=1,\dots,q.$
+Finally, the derivative with respect to the learner’s weights and the adversary’s data is given by
+$\frac{\partial^{2}f}{\partial w_{i}\partial X_{jk}}(w,X)=\frac{1}{m}\frac{\partial^{2}\mathcal{L}}{\partial w_{i}\partial X_{jk}}(\sigma(w,X_{i}),Z_{i})\quad i,j=1,\dots,m,\;k=1,\dots,q.$
+Let $X^{0}\in\mathbb{R}^{mq}$ be the start point of the adversary’s data and let the lower-level constraints $g:\mathbb{R}^{mq}\rightarrow(-1,1)$ be defined as in (6), where each constraint function $g_{i}(X):\mathbb{R}^{q}\rightarrow\mathbb{R},\;i=1,\dots,m$ measures the cosine similarity between the adversary’s data and its original position, as given by (8), where $\delta\in\mathbb{R}$ is the similarity threshold.
 The derivative of the constraints with respect to the classifier weights is 0.
-The derivative with respect to the adversary’s data is given by the following cases,
-$\frac{\partial g_{i}(X)}{\partial X_{jk}}=\begin{cases}\frac{X_{ik}^{0}}{\|X_{i}\|\cdot\|X_{i}^{0}\|}-d(X_{i},X_{i}^{0})\frac{X_{ik}}{\|X_{i}\|^{2}}&i=j\\ 0&i\neq j\end{cases}.$
-The second derivative with respect to the adversary’s data is given by the cases,
-$\frac{\partial g_{i}(X)}{X_{jk}X_{lc}}=\begin{cases}\frac{X_{ic}X_{ik}^{0}+X_{ik}X_{ic}^{0}}{\|X_{i}\|^{3}\|X_{i}^{0}\|}-\frac{3X_{ik}X_{ic}d(X_{i},X_{i}^{0})}{\|X_{i}\|^{4}}&i=j=l,k\neq c\\ \frac{2X_{ik}X_{ik}^{0}}{\|X_{i}\|^{3}\|X_{i}^{0}\|}-\frac{3X_{ik}^{2}d(X_{i},X_{i}^{0})}{\|X_{i}\|^{4}}+\frac{d(X_{i},X_{i}^{0})}{\|X_{i}\|^{2}}&i=j=l,k=c\\ 0&\text{otherwise}.\end{cases}$
+The derivative with respect to the adversary’s data is obtained as
+$\frac{\partial g_{i}(X)}{\partial X_{jk}}=\begin{cases}\frac{X_{ik}^{0}}{\|X_{i}\|\cdot\|X_{i}^{0}\|}-d(X_{i},X_{i}^{0})\frac{X_{ik}}{\|X_{i}\|^{2}}&i=j,\\ 0&i\neq j.\end{cases}$
+The second derivative with respect to the adversary’s data is given by the cases
+$\frac{\partial^{2}g_{i}(X)}{\partial X_{jk}\partial X_{lc}}=\begin{cases}\frac{X_{ic}X_{ik}^{0}+X_{ik}X_{ic}^{0}}{\|X_{i}\|^{3}\|X_{i}^{0}\|}-\frac{3X_{ik}X_{ic}d(X_{i},X_{i}^{0})}{\|X_{i}\|^{4}}&i=j=l,\;\,k\neq c,\\ \\ \frac{2X_{ik}X_{ik}^{0}}{\|X_{i}\|^{3}\|X_{i}^{0}\|}-\frac{3X_{ik}^{2}d(X_{i},X_{i}^{0})}{\|X_{i}\|^{4}}+\frac{d(X_{i},X_{i}^{0})}{\|X_{i}\|^{2}}&i=j=l,\;\,k=c,\\ \\ 0&\text{otherwise}.\end{cases}$